OpenStudy (anonymous):

Compute the indicated quantities when u = <2,1,-3> v = <4,0,1> w = <-2,6,5> * = dot product x = cross product (u x v)*w I got i-14j-4k for (u x v) How do I get the the cross product with w? My calculator says -106 but I got -2i-84j-20k

OpenStudy (ipwnbunnies):

Yeah, for (uxv)*w, don't follow what your calculator says. You computed the cross product. Now you're just multiplying that vector by a scalar multiple, w.

OpenStudy (anonymous):

So is my answer right? I just multiplied i *-2 = -2i 6*-14 = -84j -4*5 = -20k thus, -2i-84j-20k

OpenStudy (ipwnbunnies):

OH. Whoops, I didn't see that 'w' was a vector. There is no 'multplication' of two vectors, just dot product and cross product. I'm not sure. :3

OpenStudy (ipwnbunnies):

Maybe they meant the magnitude of vector w?

OpenStudy (anonymous):

well then what would be the dot product of, lets say r=<i-14j-4k> and w = <-2,6,5>: r*w

OpenStudy (ipwnbunnies):

Do you know how to calculator dot product?

OpenStudy (ipwnbunnies):

calculate*?

OpenStudy (anonymous):

yeah its would be like t=<x1,x2,x3> and r=<y1,y2,y3> t*r = x1y1+x2y2+x3y3

OpenStudy (ipwnbunnies):

Exactly. Same thing in this case. You found a vector from a cross product. You dot product the result with another vector.

OpenStudy (anonymous):

well that's how I got -2i-84j-20k I just want to make sure that that is correct. Here is my work: i-14j-4k dot product with w = <-2,6,5> i *-2 = -2i -14j*6 = -84j -4k*5 = -20k

OpenStudy (anonymous):

i-14j-4k comes from u x v

OpenStudy (ipwnbunnies):

WHOA. The result of a dot product is scalar, not a vector.

OpenStudy (anonymous):

so u x v should stay scalar?

OpenStudy (ipwnbunnies):

You said it. We're adding up the products of each corresponding component. Not making a new vector.

OpenStudy (ipwnbunnies):

No, r.w is scalar. The cross product yields another vector.

OpenStudy (ipwnbunnies):

I should've said the result of a dot product is scalar.

OpenStudy (anonymous):

Ok I see, so the calculator was right, its -106. I thought I would get a vector. -2-84-20=-106

OpenStudy (ipwnbunnies):

Yeah, that would be correct if they are actually looking for the dot product.

OpenStudy (anonymous):

So anytime I do a dot product, the answer will always be scalar?

OpenStudy (ipwnbunnies):

Yes m8. B)

OpenStudy (anonymous):

what about when I do a cross product? Will I always get a vector?

OpenStudy (ipwnbunnies):

Yes yes m8. B) And graphically, the result of the cross product will be perpendicular to the two vectors doing the product.

OpenStudy (anonymous):

Ok thx, that's the part I didn't know; that cross products equal vectors and dot products equal scalars

OpenStudy (ipwnbunnies):

YESH.